Describe the stages in development of symbolic algebra. Give at least one example of an equation for each of the stages

1
Expert's answer
2022-06-09T05:01:42-0400

Symbolic algebra, in which full symbolism is used. Early steps toward this can be seen in the work of several Islamic mathematicians such as Ibn al-Banna (13th-14th centuries) and al-Qalasadi (15th century), although fully symbolic algebra was developed by François Viète (16th century). Later, René Descartes (17th century) introduced the modern notation and showed that the problems occurring in geometry can be expressed and solved in terms of algebra.

Equally important as the use or lack of symbolism in algebra was the degree of the equations that were addressed. Quadratic equations played an important role in early algebra; and throughout most of history, until the early modern period, all quadratic equations were classified as belonging to one of three categories.

x2+px=q

x2=px+q

x2

Where p and q are positive. This trichotomy comes about because quadratic equations of the form x2+px+q=0, whis p and q positive, have no positive roots.

In between the rhetorical and syncopated stages of symbolic algebra, a geometric constructive algebra was developed by classical Greek and Vedic Indian mathematicians in which algebraic equations were solved through geometry. For instance, an equation of the form x2=A}x was solved by finding the side of a square of area A.